\begin{thebibliography}{10}

\bibitem{andersson2004functional}
M.~Andersson and J.~Sj{\"o}strand.
\newblock Functional calculus for non-commuting operators with real spectra via
  an iterated cauchy formula.
\newblock {\em Journal of Functional Analysis}, 210(2):341--375, 2004.

\bibitem{atkinson1972multiparameter}
F.~V. Atkinson and AB~Mingarelli.
\newblock {\em Multiparameter eigenvalue problems}, volume~1.
\newblock Academic Press New York, 1972.

\bibitem{bartholdi2020self}
L.~Bartholdi and S.~N. Sidki.
\newblock Self-similar products of groups.
\newblock {\em Groups, Geometry, and Dynamics}, 14(1):107--115, 2020.

\bibitem{bekka2008kazhdan}
B.~Bekka, P.~de~La~Harpe, and A.~Valette.
\newblock {\em Kazhdan's property (T)}.
\newblock Cambridge university press, 2008.

\bibitem{breuillard2017c}
E.~Breuillard, M.~Kalantar, M.~Kennedy, and N.~Ozawa.
\newblock C*-simplicity and the unique trace property for discrete groups.
\newblock {\em Publications math{\'e}matiques de l'IH{\'E}S}, 126(1):35--71,
  2017.

\bibitem{cade2013projective}
P.~Cade and R.~Yang.
\newblock Projective spectrum and cyclic cohomology.
\newblock {\em Journal of Functional Analysis}, 265(9):1916--1933, 2013.

\bibitem{conway2019course}
John~B Conway.
\newblock {\em A course in functional analysis}, volume~96.
\newblock Springer, 2019.

\bibitem{curto1988applications}
Ra{\'u}l~E Curto.
\newblock Applications of several complex variables to multiparameter spectral
  theory.
\newblock {\em Surveys of some recent results in operator theory}, 2:25--90,
  1988.

\bibitem{dixmier1982c}
J.~Dixmier.
\newblock {\em C*-algebras}.
\newblock North-Holland mathematical library. North-Holland, 1982.

\bibitem{dosi2009frechet}
A.~Dosi.
\newblock Frechet sheaves and taylor spectrum for supernilpotent lie algebra of
  operators.
\newblock {\em Mediterranean Journal of Mathematics}, 6:181--201, 2009.

\bibitem{douglas2018hermitian}
R.~G. Douglas and R.~Yang.
\newblock Hermitian geometry on resolvent set.
\newblock In {\em Operator Theory, Operator Algebras, and Matrix Theory}, pages
  167--183. Springer, 2018.

\bibitem{dudko2015spectra}
A.~Dudko and R.~Grigorchuk.
\newblock On spectra of koopman, groupoid and quasi-regular representations.
\newblock {\em arXiv preprint arXiv:1510.00897}, 2015.

\bibitem{dudko2017irreducibility}
A.~Dudko and R.~Grigorchuk.
\newblock {\em On irreducibility and disjointness of Koopman and quasi-regular
  representations of weakly branch groups}, volume 692.
\newblock American Mathematical Society Providence, RI, USA, 2017.

\bibitem{doi:10.1021/ed021p415.3}
S.~Glasstone.
\newblock Quantum chemistry (eyring, henry; walter, john; kimball, george e.).
\newblock {\em Journal of Chemical Education}, 21(8):415, 1944.

\bibitem{grigorchuk2006self}
R.~Grigorchuk and V.~Nekrashevych.
\newblock Self-similar groups, algebras and schur complements.
\newblock {\em arXiv preprint math/0612421}, 2006.

\bibitem{grigorchuk2017joint}
R.~Grigorchuk and R.~Yang.
\newblock Joint spectrum and the infinite dihedral group.
\newblock {\em Proceedings of the Steklov Institute of Mathematics},
  297:145--178, 2017.

\bibitem{haynsworth1968determination}
Emilie~V Haynsworth.
\newblock Determination of the inertia of a partitioned hermitian matrix.
\newblock {\em Linear algebra and its applications}, 1(1):73--81, 1968.

\bibitem{he2015projective}
W.~He and R.~Yang.
\newblock Projective spectrum and kernel bundle.
\newblock {\em Science China Mathematics}, 58:2363--2372, 2015.

\bibitem{hormander1973introduction}
L.~Hormander.
\newblock {\em An introduction to complex analysis in several variables}.
\newblock Elsevier, 1973.

\bibitem{liang2018quasinilpotent}
Y.~Liang and R.~Yang.
\newblock Quasinilpotent operators and non-euclidean metrics.
\newblock {\em Journal of Mathematical Analysis and Applications},
  468(2):939--958, 2018.

\bibitem{murphy2014c}
G.~J. Murphy.
\newblock {\em C*-algebras and operator theory}.
\newblock Academic press, 2014.

\bibitem{sleeman1978multiparameter}
B.~D. Sleeman.
\newblock Multiparameter spectral theory in hilbert space.
\newblock {\em Journal of Mathematical Analysis and Applications},
  65(3):511--530, 1978.

\bibitem{taylor1970joint}
J.~L. Taylor.
\newblock A joint spectrum for several commuting operators.
\newblock {\em Journal of Functional Analysis}, 6(2):172--191, 1970.

\bibitem{YANG200768}
R.~Yang.
\newblock Functional spectrum of contractions.
\newblock {\em Journal of Functional Analysis}, 250(1):68--85, 2007.

\bibitem{yang2009projective}
R.~Yang.
\newblock Projective spectrum in banach algebras.
\newblock {\em Journal of Topology and Analysis}, 1(03):289--306, 2009.

\bibitem{Yang2023JointSI}
R.~Yang.
\newblock Joint spectrum in amenability and self-similarity.
\newblock 2023.

\end{thebibliography}

\documentclass[12pt]{amsart}
\usepackage{geometry}   %????????
\usepackage[colorlinks,citecolor = red, linkcolor=blue,hyperindex]{hyperref}
\usepackage{euscript,eufrak,verbatim, mathrsfs}
\usepackage[psamsfonts]{amssymb}
\usepackage{bbm}
\usepackage{graphicx}
\usepackage{float}
\usepackage{float, tikz}

\usepackage[all, cmtip]{xy}

\usepackage{upref, xcolor, dsfont}
\usepackage{amsfonts,amsmath,amstext,amsbsy, amsopn,amsthm}
\usepackage{enumerate}

\usepackage{url}

\usepackage{mathtools}
\usepackage{bookmark}

\usepackage{euscript}
\usepackage{helvet}         % selects\textbf{\textbf{•}} Helvetica as sans-serif font
\usepackage{courier}        % selects Courier as typewriter font
\usepackage{type1cm}        % activate if the above 3 fonts are
\usepackage{multicol}        % used for the two-column index
\usepackage[bottom]{footmisc}% places footnotes at page bottom

\newtheorem{thm}{Theorem}[section]
\newtheorem{prop}[thm]{Proposition}

\newtheorem{definition}[thm]{Definition}

\numberwithin{equation}{section}

\begin{document}
	\title[Joint projective spectrum of $D_{\infty h}$]{Joint projective spectrum of $D_{\infty h}$}

	\author{Chen Li}
	\address{Chen Li: School of Mathematical Sciences, Fudan University, Shanghai, 200433, China.}
	\email{22110180024@m.fudan.edu.cn}

	\author
	{Kai Wang}
	\address{Kai WANG: School of Mathematical Sciences, Fudan University, Shanghai, 200433, China.}
	\email{kwang@fudan.edu.cn}

	\begin{abstract}
		We compute the joint spectrum of $D_{\infty h}$ with respect to the left regular representation, and finds two generators of the De Rham cohomology group of joint resolvent set which is induced by different central linear functionals. Through action of $D_{\infty h}$ on 4-ary trees, we get a self-similar realization of the group $C^*$ algebra of $D_{\infty h}$.    
	\end{abstract}

	\keywords{joint projective spectrum; maurer-cartan form; cohomology group; self similarity}

	\maketitle

	\section{Introduction}
	
	In classical Banach algebra theory, the Gelfand theory gives a comprehensive description of spectrum of operators. However, in the case of several Banach algebra elements, the (joint) spectral theory is more complicated. It is noteworthy that not only the commutativity of the tuple will influence the study, but also there is a distinction between algebraic and spatial joint spectra. 
	
	If the tuple $A=(A_1,A_2,\cdots,A_n)$ is a commutative tuple, i.e. $A_iA_j=A_jA_i$,$1\leq i,j\leq n$, J. L. Taylor defines the Taylor spectrum of the tuple by Koszul complex (see\cite{hormander1973introduction},\cite{taylor1970joint}). We refer the readers to \cite{curto1988applications}\cite{dosi2009frechet} for its applications on operator theory and sheaf theory.
	
	The matter becomes difficult when the tuple is non-commuting. In \cite{grigorchuk2017joint}, R. Grigorchuk and R. Yang generalized the classical defintion of spectrum, and considered the invertibility of the linear combination 
	\[A(z)=z_1A_1+z_2A_2+\cdots+z_nA_n.\]
	In fact, there has been increasing concern of the invertibility of $A(z)$ in fields of algebraic geometry, group theory, mathematical physics, PDEs and operator theory. We refer the readers to \cite{andersson2004functional}\cite{atkinson1972multiparameter}\cite{sleeman1978multiparameter}\cite{YANG200768} for more information. It yields the following notion of projective joint spectrum.
	
	\begin{definition}
		For a tuple $A=(A_1,A_2,\cdots,A_n)$ of elements in a unital Banach algebra $\mathcal{B}$, its projective joint spectrum $P(A)$ consists of $z\in \mathbb{C}^n$ such that $A(z)=z_1A_1+z_2A_2+\cdots+z_nA_n$ is not invertible in $\mathcal{B}$.
	\end{definition}
	
	Different from other notions of joint spectrum, for example, Taylor spectrum, the projective joint spectrum is novel in the sense of “base free". Instead of considering the invertibility of \[(A_1-z_1 I,A_2-z_2 I,\cdots, A_n-z_n I),\] we can focus on the homogeneous multiparameter pencil $A(z)$, which simplifies the study in many cases. Moreover, by homogeneousness of $A(z)$, we can consider the projective joint spectrum  $p(A)$ in complex projective space $\mathbb{P}^{n-1}=\mathbb{C}^n / \sim$ defined by $p(A)=P(A) / \sim$. In \cite{yang2009projective}, it is proved that $p(A)$ is a non-trivial compact set in $\mathbb{P}^{n-1}$ by Hartogs extension theorem.
	
	If the Banach algebra $\mathcal{B}$ is finite dimensional, matrix algebra for instance, then the projective joint spectrum is the hypersurface $\{detA(z)=0\}$. If the tuple is commutative, the projective joint spectrum will become union of hypersurfaces, which is similar to the case of Taylor spectrum.
	
	The projective resolvent set $P^c(A)=\mathbb{C}^n \text{\textbackslash}  P(A)$ and the spectrum itself has many similar properties as the single operater case. For example, it is proved that every path-connected component of $P^c(A)$ is a domain of holomorphy for general Banach algebra in \cite{he2015projective}. We also refer to the readers, for instance, \cite{cade2013projective}\cite{douglas2018hermitian}\cite{he2015projective}\cite{liang2018quasinilpotent}\cite{yang2009projective} on its applications on Hermitian metrics, hyperinvariant subspace problem and cyclic cohomology.

	Now consider a finitely generated group $G$\footnote{In this paper, all groups are discrete locally compact group unless otherwise stated.} with the generator set $S=\{g_1,g_2,\cdots,g_n\}$, and let $\rho$ be a unitary representation of $G$ on a Hilbert space $H$, which will be denoted by $(\rho, H)$. Let $C^*_{\rho}(G)$ denote the $C^*$-algebra generated by $A_i=\rho(g_i)$, $i=1,\dots,n$. The projective joint spectrum of $G$ related to $\rho$, denoted by $P(A_{\rho})$, is the projective joint spectrum of the tuple $A_\rho=(A_1,A_2,\cdots,A_n)$.
	
	 Given two representations $(\rho_1, H_1)$ and $(\rho_2, H_2)$ of $G$, they are said to be (unitarily) equivalent if there exists a unitary map $U:H_1\rightarrow H_2$ such that
	\[\rho_2(g)=U\rho_1(g)U^{-1}\qquad \forall g\in G.\]
	
	Apparently, the projective joint spectrum is invariant under equivalent representations. Moreover, it is invariant under weakly equivalent representations.  Let $(\pi, H)$ and $(\rho, K)$ be unitary representations of group $G$. We say that $\pi$ is weakly contained in $\rho$ if for every $\xi\in H$, every compact subset $Q$ of $G$ and every $\varepsilon>0$, there exists $\eta_1,\cdots,\eta_n\in K$ such that,
	\[|\langle\pi(x)\xi,\xi\rangle-\sum\limits_{i=1}^n\langle\rho(x)\eta_i,\eta_i\rangle|<\varepsilon,\quad \forall x\in Q.\]
	We shall write $\pi\prec\rho$ for this. If $\pi \prec\rho$ and $\rho\prec\pi$, we say that $\pi$ and $\rho$ are weakly equivalent and denote by $\pi\sim\rho$. Generally speaking, it is difficult to determine whether two representations are weakly equivalent from definition. However, in \cite{dixmier1982c}, it is proved that $\pi\prec\rho$ if and only if the canonical homomorphism $\rho(g)\mapsto\pi(g), g\in G$, can extend to a unital $*$-homomorphism from $C^*_{\rho}(G)$ onto $C^*_{\pi}(G)$. It implies that if $\rho(g)$ is invertible in $C^*_{\rho}(G)$, then $\pi(m)$ is invertible in $C^*_{\pi}(G)$. Moveover, $\pi\prec\rho$ implies $P(A_{\pi})\subset P(A_{\rho})$. So  $P(A_{\pi})= P(A_{\rho})$ if these two representations are weakly equivalent. 
	
	Conversely, it is a natural question whether the projective joint spectrum determines the representation up to weak equivalence. It is often not the case, but we will give some examples late in this paper. The infinite dihedral group $D_{\infty}$ is the group generated by rotations and reflections of the plane that preserves the origin. R. Yang gives detailed description on properties of joint projective spectrum of $D_{\infty}$ in \cite{Yang2023JointSI}. After introducing the concept of self-similar representation, in particular Koopman representation, it is proved that the Koopman representation and left regular representation of $D_{\infty}$ is weakly equivalent through the technique of projective joint spectrum. Koopman representation itself is widely studied in fields of ergodic theory, dynamical systems, and group representations. We refer the readers to \cite{dudko2015spectra}\cite{dudko2017irreducibility}.
	
	Throughout this article, $D_{\infty h}$ will denote the group $D_{\infty}\times \mathbb{Z}_{2}$, which is isomorphic to $\mathbb{Z}_{2} * \mathbb{Z}_{2}\times \mathbb{Z}_{2}$ and has the presentation 
	\begin{equation}\label{focusedgroup}
		D_{\infty h}=\left\langle a,t,\tau| a^2=t^2=\tau^2=1, a\tau=\tau a, t\tau=\tau t\right\rangle\text{.}
	\end{equation}
	 It can be realized as the group of rigid motions in 3-space consisting of rotations and reflections of the plane that preserves the origin with a reflection $\tau$ through the origin. Usually, the rotations are taken around the axis $Oz$ and reflections with respect to lines of the plane $Oxy$ passing through the origin, for example, see \cite{doi:10.1021/ed021p415.3} for more information and its application on electronic wave functions of molecules. %%However, although the infinite dihedral group is the union of two disjoint circles, the topological property of the projective resolvant set of $D_{\infty h}$ is totally different from that of $D_{\infty}$ as shown in section 3.

	Firstly, we will compute the projective joint spectrum of $D_{\infty h}$ with respect to its left regular representation.
	\begin{thm}
		If we define $P(R_{\lambda_{D_{\infty h}}})$ as the projective joint spectrum of \[R_{\lambda_{D_{\infty h}}}=z_0 \lambda_{D_{\infty h}}(e)+z_1\lambda_{D_{\infty h}}(a)+z_2\lambda_{D_{\infty h}}(t)+z_3\lambda_{D_{\infty h}}(\tau),\] then \[P(R_{\lambda_{D_{\infty h}}})=\bigcup\limits_{-1\leq x\leq 1}\{z\in \mathbb{C}^4:(z_0\pm z_3)^2-z_1^2-z_2^2-2z_1z_2x= 0\}.\]
	\end{thm}
	A linear functional $\phi$ on a unital Banach algebra $\mathcal{B}$ is called central if $\phi(xy)=\phi(yx)$ for any $x,y\in \mathcal{B}$. In section 3, we will concentrate on the 1-forms generated by central functionals and Maurer-Cartan form.  
	Let $Tr$ and $tr$ be the canonical tracial linear functionals on $C^*(D_{\infty h})$, $C^*(D_{\infty})$ separately, and  $\widetilde{\phi}$ be the central linear functional on $C^*(\mathbb{Z}_2)$ defined by
	\[\widetilde{\phi}(\lambda_{\mathbb{Z}_2}(e))=-1,\qquad \widetilde{\phi}(\lambda_{\mathbb{Z}_2}(\tau))=1.\] Then we will get the following theorem.
	
	\begin{thm}
		$\widetilde{\phi}\otimes tr$ induces a different element besides $Tr(\omega_R(z))$ in the cohomology group $H^1_{de}(P^c(R_{\lambda_{D_{\infty h}}}),\mathbb{Z})$.
	\end{thm}
	
	Section 4 will be devoted to give some examples to the self-similar problem raised in \cite{grigorchuk2017joint}. Finally,  We will discover relations between the Koopman representation of self similar groups and left regular representations.
	%  Koopman representation is widely studied in fields of ergodic theory, dynamical systems, and group representations. We refer the readers to \cite{dudko2015spectra}\cite{dudko2017irreducibility}. For self similar groups, its Koopman representation is also a self similar representation. We will discover relations between the Koopman representation of self similar groups and left regular representations.
	
	\begin{thm}
		The Koopman representation $\rho$ and left regular representation $\lambda$ of $D_{\infty h}$ are weakly equivalent.

	\end{thm}

	\section{Projective Joint Spectrum of \texorpdfstring{$D_{\infty h}$}{}}
	
	In this section, we will compute the projective joint spectrum of $D_{\infty h}$ with respect to the left regular representation.
	
	For a discrete group $G$, the group algebra $\mathbb{C}[G]$ is the complex linear space generated by elements in $G$, i.e.
	\[\mathbb{C}[G]=\{f|f=\sum\limits_{g\in G}a_g g, a_g\in \mathbb{C}\}\text{.}\] It is a $*$ -algebra under the conjugate operation defined by
	\begin{equation*}
		f^*=(\sum\limits_{g\in G}a_g g	)^*=\sum\limits_{g\in G}\overline{a_g}g^{-1}\text{,}
	\end{equation*}
	where $-$ represents complex conjugation and $e$ is the identity of group. Consider a positive definite function $tr$ on $\mathbb{C}[G]$ defined by $tr(f)=a_e$. By GNS construction for groups in \cite{bekka2008kazhdan}, we can get a GNS triple $(\pi_{tr},H_{tr},e_{tr})$ 
	, where $e_{tr}=e$, the Hilbert space $H_{tr}=\overline{span}\{\pi_{tr}(g)e,g\in G\} $,   and $\pi_{tr}$ is defined by $\pi_{tr}(g_1)g_2=g_1g_2$.  Let $U: H_{tr}\rightarrow l^2(G)$ be the unitary map defined by
	\[U(g)=\delta_g \quad \forall g\in G,\]
	where $\delta_g$ is the function that takes value 1 at $g$ and 0 otherwise. The left regular representation  of group $G$ on $l^2(G)$, denoted by $\lambda_G$, is defined by \[\lambda_G(s)g(t)=g(s^{-1}t) \quad \forall s \in G, g \in l^2(G).\] It is well known that $\pi_{tr}$ is unitarily equivalent to the left regular representation $\lambda_G$ of group $G$ via the map $U$. 
	
%	Recall that the reduced $C^*$-algebra  $C^*_r(G)$ is the norm closure of $\lambda_G(\mathbb{C}[G])$ in $B(l^2(G))$. In particular, if $G$ is amenable, then $C_r^*(G)$ is isomorphic to the full group $C^*$-algebra $C^*(G)$.

	By the presentation of group in \ref{focusedgroup}, $D_{\infty h}$ consists of elements that have the form of $(at)^k$, $t(at)^k$, $\tau (at)^k$, $\tau t(at)^k$. From the GNS construction mentioned above, the Hilbert space $H_{tr}$ can be decomposed as $H_{tr}=H\oplus tH\oplus \tau H\oplus \tau tH$, where
	\begin{equation*}
		H=\{f=\sum\limits_{j=-\infty}\limits^{\infty}a_j(at)^k:\sum\limits_{j=-\infty}\limits^{\infty}\left| a_j\right|^2< \infty\}\text{.}
	\end{equation*}
    Multiplication by $at$ on the Hilbert space $H$ will be denoted by $T$ in the sequel.
	By the unitary map $V\left((at)^k\right)=e^{ik\theta}$, $H$ is isomorphic to $L^2(\mathbb{T},\frac{1}{2\pi}d\theta)$ and $T$ is unitarily equivalent to the bilateral shift operator.

	Let \[R_{\lambda_{D_{\infty h}}}(z)=z_0 \lambda_{D_{\infty h}}(e)+z_1\lambda_{D_{\infty h}}(a)+z_2\lambda_{D_{\infty h}}(t)+z_3\lambda_{D_{\infty h}}(\tau)\]
	be the complex linear combination of the tuple $(\lambda_{D_{\infty h}}(e),\lambda_{D_{\infty h}}(a),\lambda_{D_{\infty h}}(t),\lambda_{D_{\infty h}}(\tau))$. Since projective joint spectrum is invariant up to unitary equivalence, we will omit the unitary operators in the following if there is no confusion.
	
	\begin{thm}\label{joint-spec}
		If we define $P(R_{\lambda_{D_{\infty h}}})$ as the projective joint spectrum of $D_{\infty h}$ related to the left regular representation, then \[P(R_{\lambda_{D_{\infty h}}})=\bigcup\limits_{-1\leq x\leq 1}\{z\in \mathbb{C}^4:(z_0\pm z_3)^2-z_1^2-z_2^2-2z_1z_2x= 0\}.\]
	\end{thm}
	
	\begin{proof}
	%	Using the spectral theorem for normal operators in \cite{conway2019course}, we can write  $T=\int_{\mathbb{T}}\lambda dE(\lambda)$, where $dE(\lambda)$ is the spectral measure of $T$.
		
		On the orthogonal direct sum $H_{tr}=H\oplus tH\oplus \tau H\oplus \tau tH$ and by the unitary map 
		$W: H\oplus tH\oplus \tau H\oplus \tau tH\rightarrow H\oplus H\oplus H\oplus H$	 defined by
		
		\begin{equation*}
			W=\begin{pmatrix}
				&I_{H}	&0  &0  &0\\
				&0  &I_{H}t	&0  &0\\
				&0  &0  &I_{H}\tau &0\\
				&0  &0  &0  &\tau t\\
			\end{pmatrix},
		\end{equation*}
		we can easily compute that
		\begin{equation}\label{matrix-representation}
			R_{\lambda_{D_{\infty h}}}(z)\overset{W}{\backsimeq}
			\begin{pmatrix}
				z_0  &z_1T+z_2  &z_3  &0\\
				z_1T^*+z_2  &z_0  &0  &z_3\\
				z_3  &0  &z_0  &z_1T+z_2\\
				0  &z_3  &z_1T^*+z_2  &z_0
			\end{pmatrix},
		\end{equation}
		where $A\overset{W}{\backsimeq} B$ means $A$ and $B$ is unitarily equivalent under $W$.
		We divide the argument into two case.
		
		\noindent Case \uppercase\expandafter{\romannumeral1}
		
		If $z_3=0$,then
		\begin{equation*}
			R_{\lambda_{D_{\infty h}}}(z)\overset{W}{\backsimeq}
			\begin{pmatrix}
				z_0  &z_1T+z_2  &0  &0\\
				z_1T^*+z_2  &z_0  &0  &0\\
				0  &0  &z_0  &z_1T+z_2\\
				0  &0  &z_1T^*+z_2  &z_0
			\end{pmatrix}.
		\end{equation*}
		Therefore, it is invertible if and only if $\begin{pmatrix}
			z_0 &z_1T+z_2\\
			z_1T^*+z_2 &z_0\\
		\end{pmatrix}$ is invertible. In this case the projective joint spectrum $P(R_{\lambda_{D_{\infty h}}})$ equals \[P(R_{\lambda_{ D_{\infty}}})=\bigcup\limits_{-1\leq x\leq 1}\{z\in \mathbb{C}^3:z_0^2-z_1^2-z_2^2-2z_1z_2x= 0\}\] from the results in \cite[Theorem 1.1]{grigorchuk2017joint}. 
		~\\
		
		\noindent Case \uppercase\expandafter{\romannumeral2}

		If $z_3\ne 0$,  $\begin{pmatrix}
			z_0 &z_1T+z_2\\
			z_1T^*+z_2 &z_0\\
		\end{pmatrix}$ will be denoted by $R_{\lambda_{D_\infty}}(\tilde{z})$ for simplicity, where $\tilde{z}=(z_0,z_1,z_2)$. Multiplying $\begin{pmatrix}
			0  &I\\
			I  &0\\
		\end{pmatrix}$ on the right,  $R_{\lambda_{ D_{\infty h}}}(z)$ turns into $\begin{pmatrix}
			z_3  &R_{\lambda_{D_\infty}}(\tilde{z})\\
			R_{\lambda_{D_\infty}}(\tilde{z})  &z_3
		\end{pmatrix}$. By the Schur complement trick\cite{haynsworth1968determination}, $R_{\lambda_{ D_{\infty h}}}(z)$ is invertible if and only if $z_3^2-R_{\lambda_{D_\infty}}^2(\tilde{z})$ is invertible.
	
	Using the spectral theorem for normal operators in \cite[Theorem 9.2.2]{conway2019course}, we can write  \begin{equation}\label{spec-reso}
		T=\int_{\mathbb{T}}\lambda dE(\lambda),
	\end{equation} where $dE(\lambda)$ is the spectral measure of $T$. 
	Note that
	\begin{eqnarray*}
	R_{\lambda_{D_\infty}}^2(\tilde{z})-z_3^2&=&\begin{pmatrix}
													z_0-z_3 &z_1T+z_2\\
													z_1T^*+z_2 &z_0-z_3
												\end{pmatrix}\begin{pmatrix}
												z_0+z_3 &z_1T+z_2\\
												z_1T^*+z_2 &z_0+z_3
											\end{pmatrix}.
	\end{eqnarray*}
This follows that  $z_3^2-R_{\lambda_{D_\infty}}^2(\tilde{z})$  is invertible if and only
 if 
 		\begin{equation*}
 			((z_0-z_3)^2-(z_1T+z_2)(z_1T^*+z_2))((z_0+z_3)^2-(z_1T+z_2)(z_1T^*+z_2))
 		\end{equation*}
 is invertible,	or more precisely, by (\ref{spec-reso})
 		\begin{equation*}
			((z_0-z_3)^2-z_1^2-z_2^2-z_1z_2(\lambda+\overline{\lambda}))((z_0+z_3)^2-z_1^2-z_2^2-z_1z_2(\lambda+\overline{\lambda}))\ne 0
		\end{equation*}
		for  each $ \lambda \in \mathbb{T}$. Letting $\lambda=e^{i\theta}$ for $ \theta \in \mathbb{T}$, $x=cos\theta$, the condition will become
		\begin{equation*}
			((z_0-z_3)^2-z_1^2-z_2^2-2z_1z_2x)((z_0+z_3)^2-z_1^2-z_2^2-2z_1z_2x)\ne 0.
		\end{equation*}
		
		In conclusion, $P^c(R_{\lambda_{D_{\infty h}}})=\bigcap\limits_{-1\leq x\leq 1}\{z\in \mathbb{C}^4:(z_0\pm z_3)^2-z_1^2-z_2^2-2z_1z_2x\ne 0\}$, and the theorem is proved by taking complement.
		
	\end{proof}
%	\begin{remark}\label{abbr-remark}
	%	\renewcommand{\qedsymbol}{}
	%	\textnormal{   Since $((z_0-z_3)^2-z_1^2-z_2^2-2z_1z_2x)$ and $((z_0+z_3)^2-z_1^2-z_2^2-2z_1z_2x)$ will frequently be used in the following sections, we will denote them by $G_x^{-}(z)$ and $G_x^{+}(z)$ separately, and $G_{\theta}^{-}(z)$, $G_{\theta}^{+}(z)$ if $x=cos\theta$.}

%	\end{remark}
	 In the following article, we will denote 
	 \begin{equation}\label{suoxie}
	 	\begin{aligned}
	 		G_x^{-}(z)&=&(z_0-z_3)^2-z_1^2-z_2^2-2z_1z_2x,\\
	 		G_x^{+}(z)&=&(z_0+z_3)^2-z_1^2-z_2^2-2z_1z_2x
	 	\end{aligned}
	 \end{equation}
	  and $G_{\theta}^{-}(z)$, $G_{\theta}^{+}(z)$ if $x=cos\theta$.
	
	\section{Trace of Maurer-Cartan Form and De Rham Cohomology group}
	%The Maurer-Cartan form is a differential 1-form on a Lie group $G$ that carries some infinitesimal information about the structure of $G$. In particular, if $G$ can be embedded in $GL(n)$ by a matrix valued mapping $g\mapsto (g_{ij})$, then the Maurer-Cartan form $\omega$ can be written as $\omega_g=g^{-1}dg$. 
	By the defintion in \cite{yang2009projective},
	for a tuple $A=(A_1,A_2,\dots,A_n)$, the Maurer-Cartan form $\omega_A$ is a operator valued 1-form defined as 
	\begin{equation*}
		\omega_A(z)=A^{-1}(z)dA(z)=\sum\limits_{i=1}^n A(z)^{-1}A_idz_i \qquad \forall z\in P^c(A).
	\end{equation*}
	
	A linear functional $\phi$ on a unital Banach algebra $\mathcal{B}$ is called central if $\phi(xy)=\phi(yx)$ for any $ x,y\in \mathcal{B}$. In \cite[Theorem 3.2]{yang2009projective}, it is proved that if $\phi$ is central and $\phi(I)\ne 0$, then $\phi(\omega_A)$ is a non-trivial element in De Rham cohomology group $H^1_{de}(P^c(A),\mathbb{C})$. This section will be devoted to 1-forms induced by different central linear functionals.
	
	For a discrete group $G$, it is well known that its reduced group $C^*$-algebra $C_r^*(G)$ admits a canonical tracial state  
	\begin{equation}\label{Ctrformula}
		tr(a)=\left<a\delta_e,\delta_e\right> \, \forall a \in C_r^*(G).
	\end{equation}
	In \cite[Corollary 4.3]{breuillard2017c}, it is proved that reduced group $C^*$-algebra $C_r^*(G)$ has only one tracial state if and only if $G$ is amenable or none of its normal subgroup is amenable. Since $D_{\infty h}$ itself is amenable, so the canonical trace $Tr $ is the unique tracial state on $C_r^*(D_{\infty h})$.
	
	Since $D_{\infty}=\mathbb{Z}\rtimes\mathbb{Z}_2$ and $D_{\infty h}=D_{\infty}\times \mathbb{Z}_2$, the inclusion map
	\[\mathbb{Z}\hookrightarrow D_{\infty}\hookrightarrow D_{\infty h}\] will induce the inclusion map of group $C^*$ algebras. Based on this observation, we have the following proposition.
	
	\begin{prop}
		The cannonical traces on $C^*_r(\mathbb{Z}), C^*_r(D_{\infty}), C^*_r(D_{\infty h})$ coincide by restriction.
	\end{prop}
	
	\begin{proof}
		Firstly, denote the cannonical traces on these $C^*$ algebras be $tr_{\mathbb{Z}}$, $tr$ and $Tr$. 
		
		By the form of elements in $D_{\infty}$, $l^2(D_\infty)=l^2(\mathbb{Z})\oplus l^2(t\mathbb{Z})$. This follows that any $a\in C^*_r(\mathbb{Z})$ can be treated as $a\oplus 0\in C^*_r(D_{\infty})$. Thus
		\[tr_\mathbb{Z}(a)=\left<a\delta_{e_\mathbb{Z}},\delta_{e_\mathbb{Z}}\right>=\left<(a\oplus 0)\delta_{e_{D_{\infty}}},\delta_{e_{D_{\infty}}}\right>=tr(a\oplus 0),\]
		which leads to $tr|_{\mathbb{Z}}=tr_{\mathbb{Z}}$.
		
		Since \[C^*_r(D_{\infty h})=\overline{C^*_r(\mathbb{Z})\otimes C^*_r(D_{\infty})}^{B(l^2(D_{\infty h}))}\]
		and $\delta_{e_{D_{\infty h}}}=\delta_{e_\mathbb{Z}}\otimes \delta_{e_{D_{\infty}}}$, for any $a\in C^*_r(D_{\infty})$,
		\[tr(a)=\left<a\delta_{e_{D_{\infty}}},\delta_{e_{D_{\infty}}}\right>=\left<a\delta_{e_{D_{\infty}}},\delta_{e_{D_{\infty}}}\right>\left<\delta_{e_\mathbb{Z}},\delta_{e_\mathbb{Z}}\right>=\left<(a\otimes 1)\delta_{e_{D_{\infty h}}},\delta_{e_{D_{\infty h}}}\right>=Tr(a\otimes 1).\]
		Therefore, $Tr|_{D_{\infty}}=tr$.

	\end{proof}
	By this proposition, we will not distinguish the cannonical traces on such groups, and denote them by $tr$.

	Noting that Group Von Neumann algebra $L(G)$ is closure of $\lambda_G(\mathbb{C}[G])$ with respect to the weak topology in the Hilbert space, thus formula (\ref{Ctrformula}) can be natrually extended to the group Von Neumann algebra. In the previous section, we get the matrix representation of $R_{\lambda_{D_{\infty h}}}$ by \eqref{matrix-representation}. In order to compute trace of the Maurer-Cartan form, we can cannonically define the extended trace $ \widetilde{tr}$ on $4\times 4$ matrices with $L(\mathbb{Z})$ entries by
	\begin{equation}\label{trace-formula}
		\widetilde{tr}((a_{ij})_{i,j=1}^4):=\frac{1}{4}tr(\sum\limits_{i=1}^4 a_{ii}).
	\end{equation}
and $\sim$ will be omitted if there is no confusion.

	By the observation in \cite{grigorchuk2017joint}, we have
	\begin{equation}\label{operator-resolution}
		tr(dE(\lambda))=tr(dE(e^{i\theta}))=dtr(E(e^{i\theta}))=\frac{1}{2\pi}d\theta,
	\end{equation} where  $E(\lambda)$ is the spectral resolution in (\ref{spec-reso}). 

For $D_{\infty h}$, the Maurer-Cartan form is
\begin{equation*}
\omega_R(z)=R^{-1}(z)dR(z)=R^{-1}(z)(dz_0+adz_1+tdz_2+\tau dz_3),
\end{equation*} where $R(z)$ is defined in (\ref{matrix-representation}). The following proposition will focus on calculating $tr(\omega_R(z))$.
	
	\begin{prop}\label{traceM}
		The trace of Maurer-Cartan form $\omega_R(z)$ is 
		\[tr(\omega_{R}(z))=\displaystyle d\bigg(\frac{1}{8\pi}\int_0^{2\pi}log((z_0-z_3)^2-z_1^2-z_2^2-2z_1z_2cos\theta)((z_0+z_3)^2-z_1^2-z_2^2-2z_1z_2cos\theta))d\theta\bigg)\]
		for $z\in P^c(R_{\lambda_{D_{\infty h}}})$.
	\end{prop}
	
	\begin{proof}
		
		As in the proof of Theorem \ref{joint-spec}, the argument will be divided into two parts.
		
		% But firstly, we need to give a specific form of $R^{-1}(z)$.
		
		\noindent Case \uppercase\expandafter{\romannumeral1}
		
		If $z_3=0$ , then $\omega_R(z)=(z_0+z_1 a+z_2 t)^{-1}(dz_0+adz_1+tdz_2)$. In this case, we can directly use the results in \cite[Propsition 3.2]{grigorchuk2017joint} to get
		\[tr(\omega_R(z))=d\bigg(\frac{1}{4\pi}\int_0^{2\pi}log(z_0^2-z_1^2-z_2^2-2z_1z_2cos\theta)d\theta\bigg),\]
		where $\displaystyle d$ stands for 
		\begin{equation*}
			\frac{\partial}{\partial z_0}dz_0+\frac{\partial}{\partial z_1}dz_1+\frac{\partial}{\partial z_2}dz_2.
		\end{equation*}
		\noindent Case \uppercase\expandafter{\romannumeral2}
		
		If $z_3\ne 0$, we need to compute $tr(R^{-1}(z))$, $tr(R^{-1}(z)a)$, $tr(R^{-1}(z)t)$, and $tr(R^{-1}(z)\tau)$ separately. By Schur complement as in the proof of Theorem \ref{joint-spec}, we get that
		\begin{equation*}
			R^{-1}(z)=\begin{pmatrix}
				-z_3^{-1}R_{\lambda_{D_\infty}}(\tilde{z})K(z)^{-1} &z_3^{-1}+z_3^{-1}R_{\lambda_{D_\infty}}(\tilde{z})K(z)^{-1}R_{\lambda_{D_\infty}}(\tilde{z})z_3^{-1}  \\
				K(z)^{-1}
				&-K(z)^{-1}R_{\lambda_{D_\infty}}(\tilde{z})z_3^{-1} 
			\end{pmatrix}
		\end{equation*}
		where $K(z)=z_3^{-1}(z_3^2-R_{\lambda_{D_\infty}}^2(\tilde{z}))$,  $R_{\lambda_{D_\infty}}(\tilde{z})=\begin{pmatrix}
			z_0 &z_1T+z_2\\
			z_1T^*+z_2 &z_0\\
		\end{pmatrix}$, and $\tilde{z}=(z_0,z_1,z_2)$.
		
		Before moving on the computation, we should take a close look at the relation between $z_0$ and $z_3$ to give a straightforward expression of $K(z)^{-1}$. It cannot hold that $z_0-z_3=0$ and $z_0+z_3=0$ simultaneously, which will contradict the precondition that $z_3\ne 0$. Thus we have three cases to discuss based on the value of $z_0-z_3$ and $z_0+z_3$.
		
		Case $(\romannumeral1)$: if $z_0=z_3$, so $z_0\ne 0$ automatically.
		A direct computation shows that \begin{eqnarray*}
			(R_{\lambda_{D_\infty}}^2(\tilde{z})-z_3^2)^{-1}&=&\begin{pmatrix}
				0  &z_1T+z_2\\
				z_1T^*+z_2  &0
			\end{pmatrix}^{-1}\begin{pmatrix}
				2z_0  &z_1T+z_2\\
				z_1T^*+z_2  &2z_0
			\end{pmatrix}^{-1}\\
			&=&\begin{pmatrix}
				0  &(z_1T^*+z_2)^{-1}\\
				(z_1T+z_2)^{-1}  &0
			\end{pmatrix}\begin{pmatrix}
				R^{11}  &R^{12}\\
				R^{21}  &R^{22}
			\end{pmatrix},
		\end{eqnarray*}
		where
		\begin{eqnarray*}
			R^{11} &=& \frac{1}{2}z_0^{-1}+\frac{1}{4}z_0^{-2}(z_1T+z_2)(2z_0-\frac{1}{2}z_0^{-1}(z_1T^*+z_2)(z_1T+z_2))^{-1}(z_1T^*+z_2),\\
			R^{12} &=& -\frac{1}{2}z_0^{-1}(z_1T+z_2)(2z_0-\frac{1}{2}z_0^{-1}(z_1T^*+z_2)(z_1T+z_2))^{-1},\\
			R^{21} &=& -(2z_0-\frac{1}{2}z_0^{-1}(z_1T^*+z_2)(z_1T+z_2))^{-1}(z_1T^*+z_2)\frac{1}{2}z_0^{-1},
		\end{eqnarray*}
		and
	\noindent		\begin{equation*}
	R^{22} \enspace=\enspace (2z_0-\frac{1}{2}z_0^{-1}(z_1T^*+z_2)(z_1T+z_2))^{-1}.\quad\qquad\qquad\qquad\qquad\qquad\qquad\qquad
		\end{equation*}
		Therefore, \begin{eqnarray*}
			tr(R^{-1}(z))&=&\frac{1}{2}tr\bigg((R_{\lambda_{D_{\infty}}}(\widetilde{z})^2-z_0^2)^{-1}R_{\lambda_{D_{\infty}}}(\widetilde{z})\bigg)\\
			&=&\frac{1}{2}tr(R^{11}+R^{12}z_0(z_1T+z_2)^{-1}+R^{21}z_0(z_1T^*+z_2)^{-1}+R^{22})\\
			&=&-\frac{1}{2}\int_{\mathbb{T}}\frac{2z_0tr(dE(\lambda))}{4z_0^2-z_1^2-z_2^2-z_1z_2(\lambda+\overline{\lambda})}\\
			&=&-\frac{1}{\pi}\int_0^{2\pi}\frac{z_0d\theta}{4z_0^2-z_1^2-z_2^2-2z_1z_2\cos\theta}.
		\end{eqnarray*}
		Using similar argument, we have that
		\begin{eqnarray*}
			tr(R^{-1}(z)a)&=&\frac{1}{2\pi}\int_0^{2\pi}\frac{(z_1+z_2cos\theta)(2z_0^2-z_1^2-z_2^2-2z_1z_2cos\theta)}{(4z_0^2-z_1^2-z_2^2-2z_1z_2cos\theta)(z_1^2+z_2^2+2z_1z_2cos\theta)}d\theta\\
			tr(R^{-1}(z)t)&=&\frac{1}{2\pi}\int_0^{2\pi}\frac{(z_1+z_2cos\theta)(2z_0^2-z_1^2-z_2^2-2z_1z_2cos\theta)}{(4z_0^2-z_1^2-z_2^2-2z_1z_2cos\theta)(z_1^2+z_2^2+2z_1z_2cos\theta)}d\theta,
		\end{eqnarray*}
		and
		\begin{equation*}
			tr(R^{-1}(z)\tau)\enspace=\enspace-\frac{1}{\pi}\int_0^{2\pi}\frac{z_0d\theta}{4z_0^2-z_1^2-z_2^2-2z_1z_2\cos\theta}d\theta.\qquad\qquad\qquad\qquad\enspace\quad
		\end{equation*}
		
		Case $(\romannumeral2)$:  $z_0+z_3=0$. In this condition, the result will be the same as case $(\romannumeral1)$.
		
		Case $(\romannumeral3)$: $z_0\ne z_3$ and $z_0\ne -z_3$. Let $G_{T}^-(z)=(z_0-z_3)^2-(z_1T+z_2)(z_1T^*+z_2)$ \\and $G_{T}^+(z)=(z_0+z_3)^2-(z_1T+z_2)(z_1T^*+z_2)$.
		
		Thus
		 \begin{eqnarray*}
			K(z)^{-1}&=&-z_3\begin{pmatrix}z_0+z_3   &z_1T+z_2\\
				z_1T^*+z_2 &z_0+z_3
			\end{pmatrix}^{-1}\begin{pmatrix}
				z_0-z_3   &z_1T+z_2\\
				z_1T^*+z_2 &z_0-z_3
			\end{pmatrix}^{-1}\\
			&=&-z_3\begin{pmatrix}
				R_{11}  &R_{12}\\
				R_{21}  &R_{22}
			\end{pmatrix},
		\end{eqnarray*}
		where
		\begin{eqnarray*}    %这地方长的离谱，之后再想想怎么修改%
			&R_{11}&=(z_0^2-z_3^2)^{-1}\bigg(I+(z_1T+z_2)G_{T}^+(z)^{-1}(z_1T^*+z_2)\bigg)\bigg(I+(z_1T
			+z_2)G_{T}^-(z)^{-1}(z_1T^*\\
			&&+z_2)\bigg)+
			(z_1T+z_2)G_{T}^+(z)^{-1}G_T^-(z)^{-1}(z_1T^*+z_2),\\
			&R_{12}&=-(z_0+z_3)^{-1}\bigg(I+(z_1T+z_2)G_T^+(z)^{-1}(z_1T^*+z_2)\bigg)(z_1T+z_2)G_T^-(z)^{-1}-(z_0-z_3)\\&&(z_1T+z_2)G_T^+(z)^{-1}(z_1T+z_2)G_T^-(z)^{-1},\\
			&R_{21}&=-(z_0-z_3)^{-1}G_T^+(z)^{-1}(z_1T^*+z_2)\bigg(I+(z_1T+z_2)G_T^-(z)^{-1}(z_1T^*+z_2)\bigg)-(z_0+z_3)\\&&G_T^+(z)^{-1}G_T^-(z)^{-1}(z_1T^*+z_2),
		\end{eqnarray*}
	and
	\begin{equation*}
R_{22}\enspace=G_T^+(z)^{-1}(z_1T^*+z_2)(z_1T+z_2)G_T^-(z)^{-1}+(z_0^2-z_3^2)G_T^+(z)^{-1}G_T^-(z)^{-1}.\qquad\qquad\;
	\end{equation*}
		
		Hence, by \eqref{trace-formula}and \eqref{operator-resolution}, we have
		\begin{eqnarray*}
			tr(R^{-1}(z))&=&\frac{1}{2}tr(-K(z)^{-1}R_{\lambda_{D_{\infty}}}(\widetilde{z})z_3^{-1})\\
			&=&\frac{1}{2}tr(z_0R_{11}+R_{12}(z_1T^*+z_2)+R_{21}(z_1T+z_2)+z_0R_{22})\\
			&=&\frac{1}{2}\int_{\mathbb{T}}\frac{2z_0(z_0^2-z_3^2)-2z_0(z_1\lambda+ z_2)(z_1\overline{\lambda}+z_2)}{G_1(z)G_2(z)}tr(dE(\lambda))\\
			&=&\frac{1}{2\pi}\int_0^{2\pi}\frac{z_0(z_0^2-z_3^2-z_1^2-z_2^2-2z_1z_2cos\theta)}{G_{\theta}^{-}(z)G_{\theta}^{+}(z)}d\theta.
		\end{eqnarray*}
		%这里分母和上面的公式太长了，所以我直接把分母写成G_1(z)G_2(z),而且具体打出全部公式出来也很长。%
		
		Similarly, we have that
				\begin{eqnarray*}
			tr(R^{-1}(z)a)&=&-\frac{1}{2\pi}\int_0^{2\pi}\frac{(z_1+z_2cos\theta)(z_0^2+z_3^2-z_1^2-z_2^2-2z_1z_2cos\theta)}{G_{\theta}^{-}(z)G_{\theta}^{+}(z)}d\theta,\\
			tr(R^{-1}(z)t)&=&-\frac{1}{2\pi}\int_0^{2\pi}\frac{(z_1cos\theta+z_2)(z_0^2+z_3^2-z_1^2-z_2^2-2z_1z_2cos\theta)}{G_{\theta}^{-}(z)G_{\theta}^{+}(z)}d\theta,
		\end{eqnarray*}
		and
		
		\begin{equation*}
			tr(R^{-1}(z)\tau)\enspace=\frac{1}{2\pi}\int_0^{2\pi}\frac{z_3(z_0^2-z_3^2-z_1^2-z_2^2-2z_1z_2cos\theta)}{G_{\theta}^{-}(z)G_{\theta}^{+}(z)}d\theta.\qquad\qquad\qquad\;
		\end{equation*}
		This follows that
		\begin{align*}
			tr(\omega_{R}(z))&=\frac{1}{4}\bigg(Tr(R^{-1}(z))dz_0+Tr(R^{-1}(z)a)dz_1+Tr(R^{-1}(z)t)dz_2+Tr(R^{-1}(z)\tau)dz_3\bigg) \\
			&=\displaystyle d\bigg(\frac{1}{8\pi}\int_0^{2\pi}log(G_{\theta}^{-}(z)G_{\theta}^{+}(z))d\theta\bigg),
		\end{align*}
		where $\displaystyle d$ stands for 
		\begin{align*}
			\frac{\partial}{\partial z_0}dz_0+\frac{\partial}{\partial z_1}dz_1+\frac{\partial}{\partial z_2}dz_2+\frac{\partial}{\partial z_3}dz_3,
		\end{align*}
		and $G_{\theta}^{-}(z), G_{\theta}^{+}(z)$ are defined in (\ref{suoxie}).	\end{proof}
		At the beginning of this section, we mentioned that a central linear functional $\phi$ will induce a non-trivial closed 1-form in De Rham cohomology group of the projective resolvent set if $\phi(I)\ne 0$. In \cite[Corollary 4.4]{grigorchuk2017joint}, it is proved by De Rham duality theorem that the  $H^1_{de}(P^c(R_{\lambda_{D_\infty}}),\mathbb{Z})$ is generated by $\frac{1}{2\pi}tr(\omega_{R_{\lambda_{D_\infty}}})$.
	So it is a natural question that what is the case for $D_{\infty h}$. In fact, there are other central linear functional will affect the cohomology group.

  It is well known that
	\begin{eqnarray*}
		C^*(G\times H)&=&C^*(G)\otimes_{max} C^*(H),\\
		C^*_r(G\times H)&=& C^*_r(G)\otimes_{min} C^*_r(H).
	\end{eqnarray*} for any discrete group $G,H$.
	Moreover, if one of the group, for example $G$, is amenable, then the group $C^*$ algebra $C^*(G)$ is nuclear and $C^*(G)$ is isomorphic to $C^*_r(G)$.  Back to our question, since both $D_{\infty}$ and $\mathbb{Z}_2$ are amenable, $C^*(D_{\infty}\times \mathbb{Z}_2)$ is the closure of  algebraic tensor product of $C^*(D_{\infty})$ and $C^*(\mathbb{Z}_2)$ in $B(l^2(D_{\infty h}))$. So tensor product of two linear functionals, on the two $C^*$ algebras respectively, will define a linear functional on $C^*(D_{\infty h})$ after extension to the closure.

	Apparently, dim$C^*(\mathbb{Z}_2)=2$. This follows that every linear functional on $C^*(\mathbb{Z}_2)$ is the complex linear combination of $\phi_1$ and $\phi_2$, where
	\begin{equation*}
		\phi_1(\lambda_{\mathbb{Z}_2}(e))=1,\qquad \phi_1(\lambda_{\mathbb{Z}_2}(\tau))=0;
	\end{equation*}
	and
	\begin{equation*}
		\phi_2(\lambda_{\mathbb{Z}_2}(e))=0,\qquad \phi_2(\lambda_{\mathbb{Z}_2}(\tau))=1.
	\end{equation*}
	 In fact, $\phi_1\otimes tr$ is just the trace on $C^*(D_{\infty h})$. Now define $\widetilde{\phi}\triangleq\phi_2-\phi_1$, it acts like twist one dimension to the opposite direction. One can check that $\widetilde{\phi}$ is not a trace, since it is -1 at the identity. The next proposition will based on the following observation.
	
	Similar to the formula of $\widetilde{tr}$ in (\ref{trace-formula}), it can be easily verified that 
	\begin{equation}\label{new-trace-formula}
		(\widetilde{\phi}\otimes tr)((a_{ij})_{i,j=1}^4)=
		-\frac{1}{4}tr(a_{11}+a_{22}+a_{33}+a_{44})+
		\frac{1}{4}tr(a_{13}+a_{24}+a_{31}+a_{42})
	\end{equation}
	
	\begin{thm}
		$\widetilde{\phi}\otimes tr$ induces a different element besides $Tr(\omega_R(z))$ in the cohomology group $H^1_{de}(P^c(R_{\lambda_{D_{\infty h}}}),\mathbb{Z})$, where $Tr$ is the cannonical trace on $C^*(D_{\infty h})$.
	\end{thm}
	
	\begin{proof}
		The proof is based on calculation of $R^{-1}(z)$ and $K^{-1}(z)$. %有个比较难受的事情是这里我想把前面关于K(z)和R^{-1}(z)的计算打个序号，然后引用的，但是前面的R^{-1}实在是太长了，后面如果要有序号的话会很丑，所以这里要再想想怎么说.
		Using the formula in \eqref{new-trace-formula}, we get%需要在附录里面写一下具体计算过程？
		\begin{equation*}
			(\widetilde{\phi}\otimes tr)(R^{-1}(z))=\frac{1}{2\pi}\int_0^{2\pi}\frac{(z_3-z_0)(z_0^2-z_3^2-z_1^2-z_2^2-2z_1z_2cos\theta)}{G_{\theta}^{-}(z)G_{\theta}^{+}(z)}d\theta,
		\end{equation*}
	and
	\begin{equation*}
		(\widetilde{\phi}\otimes tr)(R^{-1}(z)a)=\frac{1}{2\pi}\int_0^{2\pi}\frac{z_1+z_2cos\theta}{G_{\theta}^{-}(z)}d\theta.\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad
	\end{equation*}

 Choosing a point $p=(1,8,4,2)$ in the projective joint resolvent set, 
		\begin{align*}
			\displaystyle\frac{\partial\bigg(\widetilde{\phi}\otimes tr(R^{-1}(z))\bigg)}{\partial z_1}\bigg|_{z=p} &= \frac{1}{2\pi}\int_{0}^{2\pi}\frac{(16+8cos\theta)(4096cos^2\theta+10624cos\theta+6841)}{(79+64cos\theta)^2(71+64cos\theta)^2}d\theta\\
			&=\frac{14872}{45045\sqrt{105}}-\frac{7896}{45045\sqrt{2145}},
		\end{align*}
		while

		\begin{align*}
		\frac{\partial\bigg(\widetilde{\phi}\otimes tr(R^{-1}(z)a)\bigg)}{\partial z_0}\bigg|_{z=p} &=\frac{1}{2\pi}\int_{0}^{2\pi}\frac{16+8cos\theta}{(79+64cos\theta)^2}d\theta\qquad\qquad\qquad\qquad\qquad\qquad
	\\
		&=\frac{752}{2145\sqrt{2145}}.\qquad\qquad\qquad\qquad\qquad\qquad
		\end{align*}
		However, $\displaystyle\frac{\partial \bigg(Tr(R^{-1}(z)a)\bigg)}{\partial z_0}=\frac{\partial \bigg(Tr(R^{-1}(z))\bigg)}{\partial z_1}=\displaystyle\frac{\partial \bigg(Tr(\omega_{R}(z))\bigg)}{\partial z_0\partial z_1}$ by Proposition \ref{traceM}. Thus $\widetilde{\phi}\otimes tr(\omega_R(z))$ can't be multiples of $Tr(\omega_R(z))$, and definitely a different element in the cohomology group.
	\end{proof}

	As a corollary, $H^1_{de}(P^c(R_{\lambda_{D_\infty}}),\mathbb{Z})$ is not singly generated.
	
	%From the proposition, we can find a new closed 1-form in the cohomology group that is not equivalent %这个能这么说吗?%
	%		to  $tr(\omega_R(z))$. Actually, if we control the value that $\widetilde{\phi}$ get at the identity, we can get more closed 1-forms, but since the dimension of $C^*(\mathbb{Z}_2)$ is only two, so we can't find more generators in this way. So it is a natural question whether there are other central functionals on $C^*(D_{\infty h})$ that will induce new generators in the infinitely generated De Rham cohomology group.

	\section{Self Similarity of \texorpdfstring{$D_{\infty h}$}{}}
	Let $\mathcal{T}=\{x_1, x_2, \dots, x_d\}$ be a finite set with $d\geq 2$.  we denote $T^{(d)}$  the set of all finite words over $\mathcal{T}$ including the empty word $\varnothing$. It is natrually the vertex set of a d-ary tree with the empty word $\varnothing$ being its root. $T^{(d)}_n$ consists of words of length n, called the nth level of the tree. It follows that
	\begin{equation*}
		T^{(d)}=\bigcup\limits_{n=0}^{\infty} T^{(d)}_n.
	\end{equation*}
	We denote by $\displaystyle\partial T^{(d)}$ the boundary of a tree. Equipped with Tychonoff topology, it consists of all infinite length words $x_1x_2\dots$,  where $x_i\in \mathcal{T}$. Two words in $T^{(d)}$ are connected by an edge if and only if they are of the form $v$ and $xv$, where $v\in T^{(d)}, x\in \mathcal{T}$.   A self-similar action of a group $G$ is a faithful action of $G$ on $T^{(d)}$  satisfying that for every $ g\in G$ and every $ x\in \mathcal{T}$,  there exists $h\in G$ and $y\in \mathcal{T}$ such that \begin{equation*}
		g(xw)=yh(w)
	\end{equation*} and $g(\varnothing)=\varnothing$ for any $ g\in G$. By induction on the level of the tree, we can see that each $T^{(d)}_n$ is $G$-invariant and the action of $G$ on $T^{(d)}_n$ is faithful. This means that $G$ is a subgroup of the permutation group $S_{d^n}$ when restriction to each $T^{(d)}_n$ level.
	
	Let $f: T^{(d)}\rightarrow T^{(d)} $ be an endomorphism of the tree $T^{(d)}$, i.e. a map that preserves the root and adjacency of the vertices. Consider a vertex $v\in T^{(d)}$ and the subtrees $vT^{(d)}$ and $f(v)T^{(d)}$. Here $vT^{(d)}$ is the  subtree rooted at $v$ with vertices set consisting of words starting with $v$. Combining these two maps, we get a map $f: vT^{(d)}\rightarrow f(v)T^{(d)}$. 
	
	By the map $vT^{(d)}\rightarrow T^{(d)}: vw\mapsto w$, the subtree $vT^{(d)}$ is isomorphic to $T^{(d)}$. The same holds for $f(v)T^{(d)}$. After identifying $vT^{(d)}$ and $f(v)T^{(d)}$ with $T^{(d)}$, we can define an endomorphism $f|_v: T^{(d)}\rightarrow T^{(d)}$, which is uniquely determined by 
	\begin{equation*}
		f(vw)=f(v)f|_v(w).
	\end{equation*} We call $f|_v$ the restriction of $f$ in $v$.  Using this definition, a group action is self similar if and only if for any $g\in G$, $g|_v\in G$ for any $v \in T^{(d)}$.
	
	By restriction, for each fixed $n\in \mathbb{N}$ and $g\in G$, we have a wreath product decomposition $g=(g|_{v_1},g|_{v_2},\dots,g|_{v_{d^n}})\sigma$, where $v_1, v_2, \dots, v_{d^n}$ are vertices of $T^{(d)}_n$ in order, $\sigma \in S_{d^n}$, and $g_i\in Aut(T_i)$, with $T_i$ being the subtree rooted at $v_i$.
	
	A group is called self similar if it has a self similar action on some  $T^{(d)}$. It is known that the infinite dihedral group $D_{\infty}$ is a self similar group by the recursive relation
	\[ a=\sigma, \qquad t=(a,t),\]
	and $\mathbb{Z}_2$ by
	\[\tau=\sigma,\quad \text{or} \quad\tau=\sigma(\tau,\tau).\]
	In \cite{bartholdi2020self}, it is shown that the direct product of two self similar groups is self similar. In particular, if $G$ and $H$ has a self similar action on $d_1$-ary tree and $d_2$-ary tree separately, then $G\times H$ will have a self similar action on $d_1d_2$-ary tree. It follows that $D_{\infty h}$ has a self similar action on the 4-ary tree $T^{(4)}$.
	
	Next, we will define the self similar representation. A $d$-similarity of an infinite dimensional  Hilbert space $H$ is an isomorphism 
	\[\psi:H\rightarrow H^d=\underbrace{H\oplus H\oplus\dots H}_d.\]
	We can define a map $T_k: H\rightarrow H$ by
	\begin{equation}\label{coordinate-map}
		T_k(\xi)=\psi^{-1}(0,0,\dots,\xi,\dots,0)
	\end{equation}
	where  $\xi$ is the $k$-th coordinate on the right side. Note that if $\psi(\xi)=(\xi_1,\dots,\xi_d)$, then $\xi=\sum\limits_{k=1}^d T_k(\xi_k)$.  Suppose group $G$ has a self similar action on $T^{(d)}$. A unitary representation $\rho$ of $G$ on $H$ is called self similar (with respect to the d-similarity $\psi$) if
	\[\rho(g)T_x=T_y\rho(h)\]
	whenever $g(xw)=yh(w)$ for any $ w\in T^{(d)}$. By \eqref{coordinate-map}, every operator $\rho(g)\in B(H^d)$ with $g\in G$ can be written as a $d\times d$ matrix
	\[\rho(g)=(A_{yx})_{x,y\in X}\]
	where
	\[A_{yx}=\begin{cases}
		\rho(g|_x) & \text{if}\quad g(x)=y,\\
		0 & \text{otherwise}.
	\end{cases}\]
	It is called the matrix recursion on $C_{\rho}^*(G)$. 
	
	Now, let $\mu$ be the uniform Bernoulli measure on $\displaystyle\partial T^{(d)}$ and $H=L^2(\displaystyle\partial T^{(d)},\mu)$. Note that the measure $\mu$ is invariant under the induced action of self similar group $G$. The Koopman representation is defined by
	\[\rho(g)f(x)=f(g^{-1}x)\]
	for any $ g\in G$ and $f\in H$. In \cite[Example 8]{grigorchuk2006self} , it is proved that Koopman representation is  a self similar representation. In this section, we will show that the corresponding Koopman representation of $D_{\infty h}$ is weakly equivalent to the left regular representation.

%	From what we mentioned in the introduction, the projective joint spectrum is invariant under weak equivalence of group representations, since it strongly relies on the canonical isomorphism between the induced $C^*$ algebras of weakly equivalent representations. However, it is difficult to determine whether two representation are weak equivalent by definition. In \cite{grigorchuk2017joint}, projective joint spectrum helps to prove the weak equivalence of left regular representation and Koopman representation $\rho$ of the infinite dihedral group, so the $C^*$ algebra separately generated by these representations are isomorphic. We call $C^*_{\rho}(G)$ is a self similar realization of the group $C^*$ algebra. In this section, we will show that the corresponding Koopman representation of $D_{\infty h}$ is weakly equivalent to the left regular representation.
	
%	We should also note that  Koopman representation is faithful, for if $\rho(g)f=f$ for $\forall g\in G, f\in L^2(\displaystyle\partial T, \mu)$ , we can choose the function which have value 1 at a specfic point in $\displaystyle\partial T$ and 0 elsewhere to get the contradiction.
	
	\begin{thm} \label{iso-space}
		Suppose $\rho_1, \rho_2$ be the Koopman representation of groups $G$ and $H$, then they yield a Koopman representation $\rho$ of $G\times H$ and
		$C_{\rho}^*(G\times H)\cong C_{\rho_1}^*(G)\otimes_{min}C_{\rho_2}^*(H).$
		
	\end{thm}
	
	\begin{proof}
		Suppose $G$ has a self similar action  on $d_1$-ary tree with alphabet $\{x_1,x_2,\cdots, x_{d_1}\}$,  $H$ on $d_2$-ary tree with alphabet $\{y_1, y_2,\cdots y_{d_2}\}$. By \cite{bartholdi2020self}, $G\times H$ will have a self-similar action on a $d_1d_2-ary$ tree with alphabet $\{z_1,z_2,\cdots,z_{d_1d_2}\}$, and a Koopman representation $\widetilde{\rho}$. In order to define $\widetilde{\rho}$, we firstly define a map 
		\[\Phi:\displaystyle\partial T^{(d_1)}\times \displaystyle\partial T^{(d_2)}\rightarrow \displaystyle\partial T^{(d_1d_2)}\]
		by the following law: for every element $(x_1^{(d_1)}x_2^{(d_1)}\dots,y_1^{(d_2)}y_2^{(d_2)}\dots) \in \displaystyle\partial T^{(d_1)}\times \displaystyle\partial T^{(d_2)}$, we can rewrite it as $\bigg((x_1^{(d_1)},x_1^{(d_2)}),(y_2^{(d_1)},y_2^{(d_2)}),\dots\bigg)$, then $\Phi$ is defined by
		\[\Phi\bigg((x_1^{(d_1)}x_2^{(d_1)}\dots,y_1^{(d_2)}y_2^{(d_2)}\dots)\bigg)\newline=\bigg(\varphi(x_1^{(d_1)},y_1^{(d_2)}),\varphi(x_2^{(d_1)},y_2^{(d_2)}),\dots\bigg),\]where for each $n\in \mathbb{N}$
		\begin{equation*}
		\varphi(x_n^{(d_1)},y_n^{(d_2)})=z_{i+(j-1)d_1}\quad  \text{if}\quad x_n^{(d_1)}=x_i,\, y_n^{(d_2)}=y_j.
		\end{equation*}
		
		In this way, we get a one-one correspondence between $\displaystyle\partial T^{(d_1)}\times \displaystyle\partial T^{(d_2)}$ and $\displaystyle\partial T^{(d_1d_2)}$. We denote $\Phi^{-1}$ by $(\Phi_1,\Phi_2)$.
		
		Then the Koopman representation $\widetilde{\rho}$ is defined by  
		\[\widetilde{\rho}(g,h)f(\omega)=f(\Phi(g^{-1}\Phi_1(\omega),h^{-1}\Phi_2(\omega)))\] for every $(g,h)\in G\times H$, $\omega\in \displaystyle\partial T^{(d_1d_2)}$, and $f\in L^2(\displaystyle\partial T^{(d_1d_2)},\widetilde{\mu})$, where $\widetilde{\mu}$ is the Bernoulli measure on $\displaystyle\partial T^{(d_1d_2)}$.
		
		After that, we claim 
		\[L^2(\displaystyle\partial T^{(d_1)}, \mu_1)\otimes L^2(\displaystyle\partial T^{(d_2)}, \mu_2)\cong L^2(\displaystyle\partial T^{(d_1)}\times \displaystyle\partial T^{(d_2)}, \mu_1\otimes \mu_2)\cong L^2(\displaystyle\partial T^{(d_1d_2)},\widetilde{\mu}),\]
		where $\mu_1,\mu_2$ are the Bernoulli measures on $\displaystyle\partial T^{(d_1)}$ and $\displaystyle\partial T^{(d_2)}$. 
		
		The first isomorphism is a direct result from tensor product of $L^2$ spaces. For the second part, by definition of boundary of a tree, functions that depend on finite length words are dense in $L^2(\displaystyle\partial T,\mu)$, i.e. functions that have the following property:
		\[f(x_1x_2\cdots x_n\cdots)=f(x_1x_2\cdots x_n),\]
		in particular, the characteristic functions in $L^2(\displaystyle\partial T,\mu)$ that depend on finite length words. Thus $\mu_1\otimes \mu_2=\widetilde{\mu}$ is a direct result after computing the expectation of such characteristic functions. As a corollary, $\widetilde{\rho}$ is just the Koopman representation yielded by $\rho_1$ and $\rho_2$ and we will denote it by $\rho$ from now on.
		
		By definition of minimal tensor product of $C^*$-algebras and faithfulness of Koopman representation,
		\begin{eqnarray*}
			C_{\rho_1}^*(G)\otimes_{min}C_{\rho_2}^*(H)&\cong& \overline{span}\{\rho_1(g)\otimes\rho_2(h): g\in G, h\in H\}\\
			&\subseteq& B\bigg(L^2(\displaystyle\partial T^{(d_1)}, \mu_1)\otimes L^2(\displaystyle\partial T^{(d_2)}, \mu_2)\bigg).\\
		\end{eqnarray*}
		From the claim, we have
		\begin{eqnarray*}
			C_{\rho_1}^*(G)\otimes_{min}C_{\rho_2}^*(H)&\cong& \overline{span}\{U\rho_1(g)\otimes\rho_2(h)U^*: g\in G, h\in H\}\\
			&\subseteq& B\bigg(L^2(\displaystyle\partial T^{(d_1d_2)},\widetilde{\mu})\bigg).\\
		\end{eqnarray*}
		Since $C_{\rho}^*(G\times H)=\overline{span}\{\rho(g,h): (g,h)\in G\times H\}\subseteq B\bigg(L^2(\displaystyle\partial T^{(d_1d_2)},\widetilde{\mu})\bigg)$, it suffices to verify $U\rho_1(g)\otimes\rho_2(h)U^*$ and $\rho(g,h)$ has the same affect on the orthogonal basis of $L^2(\displaystyle\partial T^{(d_1)}\times \displaystyle\partial T^{(d_2)}, \mu_1\times \mu_2)$, which is trivial.
	\end{proof}
	
	Based on this proposition, we can get the following theorem directly.
	
 	%\begin{cor}\label{iso-cor}
	%	Let group $G$ be a self similar amenable group, H be a self similar group. If $\rho_G\sim \lambda_G$ and $\rho_H\sim\lambda_H$, then the same will do for $G\times H$.
	%\end{cor}

	\begin{thm}
		The Koopman representation $\rho$ and left regular representation $\lambda$ of $D_{\infty h}$ are weakly equivalent.

	\end{thm}
	
	\begin{proof}
		Suppose $\rho_1,\rho_2$ denotes the Koopman representation of $D_{\infty}$ and $\mathbb{Z}_2$. Obviously, dimension of $C^*_{\rho_{2}}(\mathbb{Z}_2)$ is at least two. By the extension of the identity map on $\mathbb{C}[\mathbb{Z}_2]$, we get a surjective $*$-homomorphism from $C^*_{\rho_{2}}(\mathbb{Z}_2)$ to $C^*(\mathbb{Z}_2)$. So as a linear space, the dimension of $C^*_{\rho_{2}}(\mathbb{Z}_2)$ is no more than that of $C^*(\mathbb{Z}_2)\cong C(\mathbb{Z}_2)$. Eventually, the $*$-homomorphism is just an isomorphism.
		
		In \cite{grigorchuk2017joint}, the left regular representation of $D_{\infty}$ is weakly equivalent to its Koopman representation $\rho_1$. Since all the mentioned $C^*$ algebras are nuclear, by Kirchberg Theorem\cite{murphy2014c},
		we get a canonical isomorphism between $C^*_r(D_{\infty})\otimes_{min} C^*(\mathbb{Z}_2)$ and $C^*_{\rho_1}(D_{\infty})\otimes_{min} C_{\rho_2}^*(\mathbb{Z}_2)$, so is the case between $C_r^*(D_{\infty h})$ and $C_{\rho}^*(D_{\infty h})$ and the theorem  follows.
	\end{proof}

	\bibliographystyle{plain}
	\bibliography{cite}

\end{document}